# VariFace: Fair and Diverse Synthetic Dataset Generation for Face Recognition

### v1

@@ 1 Introduction
A decade after the breakthrough performances of DeepFace [65] and DeepID [64], deep learning remains the state-of-the-art approach for face recognition (FR) [69, 19]. Deep learning performance is limited by training dataset size [79], and the creation of large-scale FR datasets such as CASIA-WebFace [75], MS-Celeb-1M [18] and MegaFace [27] was central to the success of deep learning in FR. However, the development of these massive face datasets involves scraping face image data from the internet without permission from subjects, and all but one of the aforementioned datasets have since been retracted or decommissioned [7]. Legally, Article 9 (1) General Data Protection Regulation (GDPR) [50] in the EU prohibits the processing of biometric data for the purpose of uniquely identifying a natural person, except when the data subject has given explicit consent to the processing of those personal data for one or more specified purposes (Article 9(2)(a) GDPR). Moreover, the AI Act (AIA) in the EU came into effect on August 1st, 2024, where Article 5(1)(e) AIA prohibits the placing on the market or the use of AI systems that create or expand facial recognition databases through the untargeted scraping of facial images [51]. Together, these regulations illustrate a growing legal concern over the use of large-scale web-scraped face datasets for training face recognition models.
@@
Besides privacy concerns, real face datasets suffer from data imbalance, including limited face pose and lighting variation [37], as well as under-representation of protected characteristics such as race and gender [37, 25, 66]. This results in FR models trained on these datasets exhibiting robustness and fairness issues, with the latter raising significant legal and ethical implications [42].
@@
To address the privacy risks and biases associated with real face datasets, there is increasing interest in the development of synthetic face datasets [3, 53, 38, 2, 8, 43, 30]. Moreover, synthetic methods not only facilitate scalable dataset generation to support deep FR model training [67, 2], but also provide precise control over the composition of the dataset [38, 43, 4, 3].
@@
The most widely used synthetic data generation methods are 3D parametric [71, 2] and deep generative [67, 53, 38, 9, 8, 43, 30] models, the latter encompassing GAN [15, 63, 53, 7, 38, 9, 11, 72] and diffusion-based approaches [8, 43, 30, 24, 49, 36]. Despite the considerable progress in the development of synthetic face data, there remains a substantial gap in FR performance between models trained on real compared to synthetic data [10, 61, 13]. This is because synthetic datasets tend to amplify biases inherent in the real face datasets used for training [66, 33], while also introducing unique challenges that affect FR performance [2, 30].
@@
To mitigate the demographic biases in real datasets, several synthetic methods are able to generate datasets with balanced demographic distributions [53, 30, 43]. However, these methods rely on using either existing demographic labels or labels generated by supervised models [65, 30, 43], which exhibit poor performance on minority classes (see Fig. S2). Another issue inherent with synthetic methods is the challenge of balancing identity (ID) preservation with generating diverse images for each individual. Although ID preservation can be achieved using pretrained FR model embeddings as a conditioning signal [8, 30, 36], generating sufficient diversity to reflect ‘in-the-wild’ variation remains challenging. To generate image variations for an individual, current methods rely on extracting specific attributes such as face pose, expression, and illumination [43, 49]. However, this requires using supervised models trained on those attributes and ignores properties that cannot be easily classified but are important for FR performance. In summary, while synthetic data generation methods achieve certain aspects, such as balanced demographics and ID preservation, the issue of generating sufficient diversity across and within identities remains unsolved.
Here we propose VariFace, a synthetic face generation pipeline that achieves fair interclass variation and diverse intraclass variation.
@@
To obtain accurate demographic labels, we first leverage a pretrained CLIP encoder [54] to extract initial predictions. These labels are subsequently refined with dataset-level information by enforcing consistency within FR embedding space. The race and gender labels are used as conditioning signals for the first stage to generate a demographically-balanced dataset of face identities. Moreover, interclass diversity is further improved by applying the Vendi score [17] as a guidance loss function during sampling. To generate diverse intraclass variation, we propose Divergence Score Conditioning, a metric in the FR embedding space that enables control over the ID preservation-intraclass diversity trade-off. This avoids the need to manually specify image attributes, greatly simplifying the generation pipeline by avoiding the use of auxiliary supervised models. By conditioning with divergence score, ID, and age labels, the second stage generates diverse but ID-preserved variation across individuals.
@@ 2 Related work
Synthetic Face Generation. The applications of generating synthetic faces are extensive and include face recognition [10, 2, 8, 43, 30, 24], reconstruction [56, 57], editing [63, 41, 52, 73] and analysis [35, 71].
@@
Using 3D face scan data, Wood et al. [71] applied a face model with a library of hair, clothing, and accessory assets to render a face dataset. 3D scan data avoids the need to use real face data while maintaining the generation of well-preserved identities. However, the acquisition of 3D face scans is time-consuming and expensive due to the need for specialized scanner equipment and considerable post-processing requirements. Moreover, there remains a significant domain gap between rendered and real faces, and the variation is limited by the assets available.
@@
In contrast, deep generative methods such as GANs and diffusion models address domain gap issues by leveraging real datasets to learn the generation of realistic faces. DiscoFaceGAN [15] combines 3D priors with an adversarial learning framework to generate faces with independent control over ID, pose, expression, and illumination. Similarly, InterFaceGAN [63] enables precise control of facial attributes while preserving ID by manipulating the latent representation along semantically meaningful directions. Despite high controllability, GANs suffer from mode collapse, resulting in limited interclass and intraclass diversity. In contrast, diffusion models facilitate diverse face image generation while retaining capabilities for fine-grained, controllable face editing [41, 30].
@@
Synthetic Datasets for Face Recognition. There is a growing interest in applying synthetic datasets for training FR models, with a surge in related competitions held in recent years [61, 13, 44]. Synthetic data provides a cost-effective method to scale datasets, facilitating the training of deep FR models that require many unique faces with a diverse set of images per individual [79, 1].
@@
Based on 3D face scan data, the DigiFace-1M [2] dataset comprises over a million rendered faces. While 3DCG data avoids privacy concerns and issues with ID preservation, the domain gap and low intraclass diversity limit FR performance [55]. In contrast, deep generative models suffer from problems with ID preservation and image artifacts but offer more flexibility to generate diversity across and within individuals. SynFace [53] leverages DiscoFaceGAN for face generation and introduces ID mixup to improve interclass diversity. However, SynFace suffers from low intraclass diversity, and IDiff-Face [8] applies dropout during training to prevent overfitting to ID context and improve intraclass diversity. Similarly, to enhance intraclass diversity, DCFace [30] uses a bank of real images for style information. However, there is a risk of data leakage using real images for style transfer, and these methods may not help to improve diversity beyond the training dataset. Instead, GANDiffFace [43] uses supervised labels from pretrained models to generate individuals of diverse ages, as well as the same individual with diverse poses and expressions. While supervised attributes can further improve intraclass diversity, the limited attributes specified may not capture all the variations important for FR. The most recent methods, such as VIGFace [31], Arc2Face [49], and Vec2Face [72] focus on generating ID embeddings to synthesize diverse face images. However, these methods do not address fairness concerns, inheriting demographic biases from the real dataset (see Sec. 4.5), and there remains a considerable gap between the performance of FR models trained on these datasets compared to real datasets.
@@ 3.1 Stage 1: Fair interclass variation
The aim of the first stage of the pipeline is to generate a diverse set of face identities with a balanced representation of races and genders. To create demographic labels, we first obtain initial estimates using a pretrained CLIP model [54, 68], and subsequently refine these predictions using a pretrained FR model. With these labels, we train a conditional diffusion model [21, 47, 16, 58, 12] to learn to generate faces with control over race and gender attributes. During sampling, we apply the Vendi score [17] as a guidance loss function to generate diversity within each demographic category. Finally, we apply an automatic filtering process to retain images that are demographic-consistent and display good face image quality for use as synthetic identities in the second stage.
Label refinement with Face Recognition Consistency. Previous methods used supervised approaches such as DeepFace [65] to extract demographic labels from face images [43, 25]. Instead, we leverage a pretrained CLIP model to obtain race, gender, and age labels (see Suppl. Sec. 6). While DeepFace uses separate age, race, and gender models, CLIP benefits from a unified embedding space and offers greater flexibility in defining labels. However, both supervised methods and CLIP generate predictions independently for each face image, ignoring crucial contextual information present in the dataset. Face embeddings extracted using FR models not only contain ID information but also the space of face embeddings appears well structured with respect to demographic information [34, 33]. Therefore, we propose to leverage the structure in the FR embedding space to refine the race and gender labels.
@@
Enhancing diversity with Face Vendi Score Guidance. Generating a diverse set of face identities is important for training FR models [30]. The Vendi score (VS) is an evaluation metric that measures diversity at a dataset level and is defined as the exponential of the Shannon entropy applied to the normalized eigenvalues of a kernel similarity matrix.
@@
Here, we propose using the VS as a guidance function to improve interclass diversity. In contrast to conditional image generation [20, 5], guided image generation involves using a pretrained frozen diffusion model to control image generation [28, 6]. By specifying a guidance loss function, the sampling process can be guided to simultaneously optimize the guidance loss function and denoising objective. Concretely, for each batch of denoised images, we obtain face embeddings $E\in\mathbb{R}^{n}$ using a pretrained FR model and compute the VS loss ($\mathcal{L}_{VS}$) specifying the dataset as the batch of embeddings $B=\{E\}_{i=1}^{m}$ and similarity function as the cosine similarity $C_{S}:\mathbb{R}^{n}\times\mathbb{R}^{n}\to\mathbb{R}$ defined in Eq. 1:
@@
The Face Vendi Score Guidance algorithm for a generic diffusion sampler $S(\cdot,\cdot,\cdot)$ is summarized in Algorithm 1. Similar to Universal Guidance [6], we apply the guidance loss to the denoised image. Specifically, we use a pretrained FR model on the batch of denoised images to generate a set of face embeddings as input into $\mathcal{L}_{VS}$. The gradient of $\mathcal{L}_{VS}$ is then scaled and combined with the noise estimated by the diffusion model and used by the diffusion sampler to generate the prediction for the following timestep. By maximizing the Vendi Score loss across timesteps, the batch of faces generated is guided to become diverse with respect to face ID, therefore maximizing interclass diversity.
@@
Filtering. To create a balanced, high-quality set of face identities, we apply a two-stage filtering process to the generated images. Firstly, we obtain demographic labels using CLIP-FRC and remove images that are inconsistent with the intended label. Secondly, to remove images with extreme poses, poor lighting, or artifacts, we use CLIB-FIQA [48] to extract face image quality labels and remove images with a score below a predefined threshold.
@@ 3.2 Stage 2: Diverse intraclass variation
Divergence score conditioning. To generate diverse variation for a given individual, it is common to specify a set of attributes such as pose or lighting, either using text prompts [3], or labels derived from classifiers [43]. However, these methods are non-exhaustive and may struggle to capture variations that cannot be easily conveyed in natural language or classified. Instead, we aim to capture diversity more holistically and without the reliance on named attributes, greatly simplifying the generation pipeline by removing the need for additional models. To do so, we assign a score to each training image based on how far the image deviates from the ‘prototypical example’ for that individual and use this score as a label for training a conditional diffusion model.
@@
Filtering. There is an inherent trade-off between intraclass diversity and ID preservation [30]. While controlling the DS helps to balance the trade-off, we further enforce this by applying a filter to remove cases where the ID was not preserved. Using the pretrained FR model, we compute the cosine similarity between the embedding from the base image generated in stage 1, with the images generated in stage 2 using that ID as a condition. We then remove images where the cosine similarity between the embeddings is below a predefined threshold.

### v2

@@ 1 Introduction
A decade after the breakthrough performances of DeepFace (Taigman et al., 2014) and DeepID (Sun et al., 2014), deep learning remains the state-of-the-art approach for face recognition (FR) (Wang & Deng, 2021; Gururaj et al., 2024). Deep learning performance is limited by training dataset size (Zhu et al., 2021), and the creation of large-scale FR datasets such as CASIA-WebFace (Yi et al., 2014), MS-Celeb-1M (Guo et al., 2016) and MegaFace (Kemelmacher-Shlizerman et al., 2016) was central to the success of deep learning in FR. However, the development of these massive face datasets involves scraping face image data from the internet without permission from subjects, and all but one of the aforementioned datasets have since been retracted or decommissioned (Boutros et al., 2022). Legally, Article 9 (1) General Data Protection Regulation (GDPR) (Parliament, 2016) in the EU prohibits the processing of biometric data for the purpose of uniquely identifying a natural person, except when the data subject has given explicit consent to the processing of those personal data for one or more specified purposes (Article 9(2)(a) GDPR). Moreover, the AI Act (AIA) in the EU came into effect on August 1st, 2024, where Article 5(1)(e) AIA prohibits the placing on the market or the use of AI systems that create or expand facial recognition databases through the untargeted scraping of facial images (Parliament, 2024). Together, these regulations illustrate a growing legal concern over the use of large-scale web-scraped face datasets for training face recognition models.
@@
Besides privacy concerns, real face datasets suffer from data imbalance, including limited face pose and lighting variation (Liu et al., 2022a), as well as under-representation of protected characteristics such as race and gender (Liu et al., 2022a; Karkkainen & Joo, 2021; Thong et al., 2023). This results in FR models trained on these datasets exhibiting robustness and fairness issues, with the latter raising significant legal and ethical implications (Mehrabi et al., 2021).
@@
To address the privacy risks and biases associated with real face datasets, there is increasing interest in the development of synthetic face datasets (Baltsou et al., 2024; Qiu et al., 2021; Liu et al., 2022b; Bae et al., 2023; Boutros et al., 2023a; Melzi et al., 2023; Kim et al., 2023). Moreover, synthetic methods not only facilitate scalable dataset generation to support deep FR model training (Trigueros et al., 2021; Bae et al., 2023), but also provide precise control over the composition of the dataset (Liu et al., 2022b; Melzi et al., 2023; Banerjee et al., 2023; Baltsou et al., 2024).
@@
The most widely used synthetic data generation methods are 3D parametric (Wood et al., 2021; Bae et al., 2023) and deep generative (Trigueros et al., 2021; Qiu et al., 2021; Liu et al., 2022b; Boutros et al., 2023b; a; Melzi et al., 2023; Kim et al., 2023) models, the latter encompassing GAN (Deng et al., 2020; Shen et al., 2020; Qiu et al., 2021; Boutros et al., 2022; Liu et al., 2022b; Boutros et al., 2023b; 2024; Wu et al., 2024a) and diffusion-based approaches (Boutros et al., 2023a; Melzi et al., 2023; Kim et al., 2023; Kansy et al., 2023; Papantoniou et al., 2024; Li et al., 2024). Despite the considerable progress in the development of synthetic face data, there remains a substantial gap in FR performance between models trained on real compared to synthetic data (Boutros et al., 2023c; Shahreza et al., 2024; DeAndres-Tame et al., 2024). This is because synthetic datasets tend to amplify biases inherent in the real face datasets used for training (Thong et al., 2023; Leyva et al., 2024), while also introducing unique challenges that affect FR performance (Bae et al., 2023; Kim et al., 2023).
@@
To mitigate the demographic biases in real datasets, several synthetic methods are able to generate datasets with balanced demographic distributions (Qiu et al., 2021; Kim et al., 2023; Melzi et al., 2023). However, these methods rely on using either existing demographic labels or labels generated by supervised models (Taigman et al., 2014; Kim et al., 2023; Melzi et al., 2023), which exhibit poor performance on minority classes (see Fig. S2). Another issue inherent with synthetic methods is the challenge of balancing identity (ID) preservation with generating diverse images for each individual. Although ID preservation can be achieved using pretrained FR model embeddings as a conditioning signal (Boutros et al., 2023a; Kim et al., 2023; Li et al., 2024), generating sufficient diversity to reflect ‘in-the-wild’ variation remains challenging. To generate image variations for an individual, current methods rely on extracting specific attributes such as face pose, expression, and illumination (Melzi et al., 2023; Papantoniou et al., 2024). However, this requires using supervised models trained on those attributes and ignores properties that cannot be easily classified but are important for FR performance. In summary, while synthetic data generation methods achieve certain aspects, such as balanced demographics and ID preservation, the issue of generating sufficient diversity across and within identities remains unsolved.
@@
Here we propose VariFace, a synthetic face generation pipeline that achieves fair interclass variation and diverse intraclass variation. To obtain accurate demographic labels, we first leverage a pretrained CLIP encoder (Radford et al., 2021) to extract initial predictions. These labels are subsequently refined with dataset-level information by enforcing consistency within FR embedding space. The race and gender labels are used as conditioning signals for the first stage to generate a demographically-balanced dataset of face identities. Moreover, interclass diversity is further improved by applying the Vendi score (Friedman & Dieng, 2023) as a guidance loss function during sampling. To generate diverse intraclass variation, we propose Divergence Score Conditioning, a metric in the FR embedding space that enables control over the ID preservation-intraclass diversity trade-off. This avoids the need to manually specify image attributes, greatly simplifying the generation pipeline by avoiding the use of auxiliary supervised models. By conditioning with divergence score, ID, and age labels, the second stage generates diverse but ID-preserved variation across individuals.
@@ 2 Related work
Synthetic Face Generation. The applications of generating synthetic faces are extensive and include face recognition (Boutros et al., 2023c; Bae et al., 2023; Boutros et al., 2023a; Melzi et al., 2023; Kim et al., 2023; Kansy et al., 2023), reconstruction (Richardson et al., 2016; 2017), editing (Shen et al., 2020; Matsunaga et al., 2022; Plesh et al., 2023; Wu et al., 2024b) and analysis (Li & Deng, 2020; Wood et al., 2021).
@@
Using 3D face scan data, Wood et al. (2021) applied a face model with a library of hair, clothing, and accessory assets to render a face dataset. 3D scan data avoids the need to use real face data while maintaining the generation of well-preserved identities. However, the acquisition of 3D face scans is time-consuming and expensive due to the need for specialized scanner equipment and considerable post-processing requirements. Moreover, there remains a significant domain gap between rendered and real faces, and the variation is limited by the assets available.
@@
In contrast, deep generative methods such as GANs and diffusion models address domain gap issues by leveraging real datasets to learn the generation of realistic faces. DiscoFaceGAN (Deng et al., 2020) combines 3D priors with an adversarial learning framework to generate faces with independent control over ID, pose, expression, and illumination. Similarly, InterFaceGAN (Shen et al., 2020) enables precise control of facial attributes while preserving ID by manipulating the latent representation along semantically meaningful directions. Despite high controllability, GANs suffer from mode collapse, resulting in limited interclass and intraclass diversity. In contrast, diffusion models facilitate diverse face image generation while retaining capabilities for fine-grained, controllable face editing (Matsunaga et al., 2022; Kim et al., 2023).
@@
Synthetic Datasets for Face Recognition. There is a growing interest in applying synthetic datasets for training FR models, with a surge in related competitions held in recent years (Shahreza et al., 2024; DeAndres-Tame et al., 2024; Melzi et al., 2024). Synthetic data provides a cost-effective method to scale datasets, facilitating the training of deep FR models that require many unique faces with a diverse set of images per individual (Zhu et al., 2021; An et al., 2022).
@@
Based on 3D face scan data, the DigiFace-1M (Bae et al., 2023) dataset comprises over a million rendered faces. While 3DCG data avoids privacy concerns and issues with ID preservation, the domain gap and low intraclass diversity limit FR performance (Rahimi et al., 2024). In contrast, deep generative models suffer from problems with ID preservation and image artifacts but offer more flexibility to generate diversity across and within individuals. SynFace (Qiu et al., 2021) leverages DiscoFaceGAN for face generation and introduces ID mixup to improve interclass diversity. However, SynFace suffers from low intraclass diversity, and IDiff-Face (Boutros et al., 2023a) applies dropout during training to prevent overfitting to ID context and improve intraclass diversity. Similarly, to enhance intraclass diversity, DCFace (Kim et al., 2023) uses a bank of real images for style information. However, there is a risk of data leakage using real images for style transfer, and these methods may not help to improve diversity beyond the training dataset. Instead, GANDiffFace (Melzi et al., 2023) uses supervised labels from pretrained models to generate individuals of diverse ages, as well as the same individual with diverse poses and expressions. While supervised attributes can further improve intraclass diversity, the limited attributes specified may not capture all the variations important for FR. The most recent methods, such as VIGFace (Kim et al., 2024), Arc2Face (Papantoniou et al., 2024), and Vec2Face (Wu et al., 2024a) focus on generating ID embeddings to synthesize diverse face images. However, these methods do not address fairness concerns, inheriting demographic biases from the real dataset (see Sec. 4.5), and there remains a considerable gap between the performance of FR models trained on these datasets compared to real datasets.
@@ 3.1 Stage 1: Fair interclass variation
The aim of the first stage of the pipeline is to generate a diverse set of face identities with a balanced representation of races and genders. To create demographic labels, we first obtain initial estimates using a pretrained CLIP model (Radford et al., 2021; Wang et al., 2023), and subsequently refine these predictions using a pretrained FR model. With these labels, we train a conditional diffusion model (Ho et al., 2020; Nichol & Dhariwal, 2021; Dhariwal & Nichol, 2021; Rombach et al., 2022; Crowson et al., 2024) to learn to generate faces with control over race and gender attributes. During sampling, we apply the Vendi score (Friedman & Dieng, 2023) as a guidance loss function to generate diversity within each demographic category. Finally, we apply an automatic filtering process to retain images that are demographic-consistent and display good face image quality for use as synthetic identities in the second stage.
Label refinement with Face Recognition Consistency. Previous methods used supervised approaches such as DeepFace (Taigman et al., 2014) to extract demographic labels from face images (Melzi et al., 2023; Karkkainen & Joo, 2021). Instead, we leverage a pretrained CLIP model to obtain race, gender, and age labels (see Suppl. Sec. A). While DeepFace uses separate age, race, and gender models, CLIP benefits from a unified embedding space and offers greater flexibility in defining labels. However, both supervised methods and CLIP generate predictions independently for each face image, ignoring crucial contextual information present in the dataset. Face embeddings extracted using FR models not only contain ID information but also the space of face embeddings appears well structured with respect to demographic information (Li et al., 2023; Leyva et al., 2024). Therefore, we propose to leverage the structure in the FR embedding space to refine the race and gender labels.
@@
Enhancing diversity with Face Vendi Score Guidance. Generating a diverse set of face identities is important for training FR models (Kim et al., 2023). The Vendi score (VS) is an evaluation metric that measures diversity at a dataset level and is defined as the exponential of the Shannon entropy applied to the normalized eigenvalues of a kernel similarity matrix.
@@
Here, we propose using the VS as a guidance function to improve interclass diversity. In contrast to conditional image generation (Ho & Salimans, 2022; Bansal et al., 2023), guided image generation involves using a pretrained frozen diffusion model to control image generation (Kim et al., 2022a; Bansal et al., 2024). By specifying a guidance loss function, the sampling process can be guided to simultaneously optimize the guidance loss function and denoising objective. Concretely, for each batch of denoised images, we obtain face embeddings $E\in\mathbb{R}^{n}$ using a pretrained FR model and compute the VS loss ($\mathcal{L}_{VS}$) specifying the dataset as the batch of embeddings $B=\{E\}_{i=1}^{m}$ and similarity function as the cosine similarity $C_{S}:\mathbb{R}^{n}\times\mathbb{R}^{n}\to\mathbb{R}$ defined in Eq. 1:
@@
The Face Vendi Score Guidance algorithm for a generic diffusion sampler $S(\cdot,\cdot,\cdot)$ is summarized in Algorithm 1. Similar to Universal Guidance (Bansal et al., 2024), we apply the guidance loss to the denoised image. Specifically, we use a pretrained FR model on the batch of denoised images to generate a set of face embeddings as input into $\mathcal{L}_{VS}$. The gradient of $\mathcal{L}_{VS}$ is then scaled and combined with the noise estimated by the diffusion model and used by the diffusion sampler to generate the prediction for the following timestep. By maximizing the Vendi Score loss across timesteps, the batch of faces generated is guided to become diverse with respect to face ID, therefore maximizing interclass diversity.
@@
Filtering. To create a balanced, high-quality set of face identities, we apply a two-stage filtering process to the generated images. Firstly, we obtain demographic labels using CLIP-FRC and remove images that are inconsistent with the intended label. Secondly, to remove images with extreme poses, poor lighting, or artifacts, we use CLIB-FIQA (Ou et al., 2024) to extract face image quality labels and remove images with a score below a predefined threshold.
@@ 3.2 Stage 2: Diverse intraclass variation
Divergence score conditioning. To generate diverse variation for a given individual, it is common to specify a set of attributes such as pose or lighting, either using text prompts (Baltsou et al., 2024), or labels derived from classifiers (Melzi et al., 2023). However, these methods are non-exhaustive and may struggle to capture variations that cannot be easily conveyed in natural language or classified. Instead, we aim to capture diversity more holistically and without the reliance on named attributes, greatly simplifying the generation pipeline by removing the need for additional models. To do so, we assign a score to each training image based on how far the image deviates from the ‘prototypical example’ for that individual and use this score as a label for training a conditional diffusion model.
@@
Filtering. There is an inherent trade-off between intraclass diversity and ID preservation (Kim et al., 2023). While controlling the DS helps to balance the trade-off, we further enforce this by applying a filter to remove cases where the ID was not preserved. Using the pretrained FR model, we compute the cosine similarity between the embedding from the base image generated in stage 1, with the images generated in stage 2 using that ID as a condition. We then remove images where the cosine similarity between the embeddings is below a predefined threshold.